OpenStudy (anonymous):

find the height of a right circular cylinder that has a volume of 62.8in cubed and a diameter of 4in

OpenStudy (paki):

simple... use formula... h=v/pie * r^2 h= 62.8/3.14*4*4 h=1.25in

OpenStudy (accessdenied):

4in is the diameter, not the radius. If you recall, the formula for volume of a cylinder: \( V = \pi r^2 h \) We want to find h. We know V (given) and r (half the diameter). \( 62.8 = \pi \left( \dfrac{4}{2} \right)^2 h \) \( 62.8 = \pi 2^2 h = 4 \pi h \)

OpenStudy (paki):

ahh thanks... then first find radius, by the formula... r=d/2.. .then put that in the formula... it will be like, h= 62.8/3.14*2*2 h= 5 inches...

OpenStudy (anonymous):

thanks for the help. i have another question. what are the slope and the y-intercept of the line y= -2x + 6?

OpenStudy (paki):

@Vickicasa

OpenStudy (paki):

and the graph will be like....

OpenStudy (paki):

got it dear...